# A Kalman Filter for track reconstruction in very large time projection chambers

### v1

@@ 3.2 Tests and results: parameter scan sample
The PS sample was further used to test the improvement in $q/p_{\text{T}}$ resolution brought by the introduction of the “mirror rotation” technique in the CKF, compared to BKF. The fraction of the total tracks for which the mirroring technique was used $\epsilon_{\textrm{Mirror}}$ is shown in Fig. 13 as a function of $t_{\textrm{ID}}$, initial true $p_{\textrm{T}}$ and $P_{\textrm{gas}}$, with primaries and secondaries being shown separately. In general, primary particles are less likely to produce looping particles than secondaries. From Fig. 13(a) we can see that the likelihood of the particles producing looping trajectory drops significantly with mass due to the higher $\textrm{d}E/\textrm{d}x$, with electrons having $\epsilon_{\textrm{Mirror}}\sim 0.2$ and protons having $\epsilon_{\textrm{Mirror}}<0.01$. From Fig. 13(b) we can also see that the only particles that loop in the detector have low initial transverse momenta $p_{\textrm{T}}<0.3\ \textrm{GeV}/c$, which comes directly from the relationship between momentum and curvature shown in Eq. 14. Finally from Fig. 13(c) we see that while for primaries the probability of producing looping trajectory remains roughly constant in this range, for secondaries it increases significantly at lower pressures.
Figure 13: Portion of tracks in the PS sample for which the mirror rotation was applied, $\epsilon_{\textrm{Mirror}}$, as a function of (a) ALICE particle identification convention $t_{\textrm{ID}}=(0,1,2,3,4)=(\textrm{e},\mu,\pi,K,\textrm{p})$, (b) gas pressure $P_{\textrm{gas}}$ and (c) transverse momentum $p_{\textrm{T}}$. Primary and Secondary tracks are drawn separately.
@@
Figure 14: Ratios of the $q/p_{\text{T}}$ resolutions obtained using the full CKF algorithm including the mirror rotation method $\sqrt{C_{44}^{\textrm{BKF}}}$, over the reconstruction without mirror rotation $\sqrt{C_{44}^{\text{BKF}}}$. The plots were produced for the whole PS sample. The histograms are color coded according to the gas pressure $P_{\textrm{gas}}$. Plot (a) contains only electrons, plot (b) contains pions and muons and plot (c) contains kaons and protons. Only tracks with a minimum of 30 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
Figure 15: Similar plots to Fig. 14. The plot contains all particle types in the PS sample. The histograms are color coded according to the number of points in the track $N$.
The difference between the results is shown to be less dramatic in more pressurized environments, where particles tend to be absorbed sooner and tracks are generally shorter. This trend is confirmed looking at the results obtained for the muons and pions sample in Fig.. 14(b), for which $\textrm{d}E/\textrm{d}x$ will on average be higher due to their higher masses: the improvement in this case is by $\sim 20\%$ for $p_{\text{T}}<150$ MeV/$c$ with peaks of up to $\sim 50$% for the lowest momentum tracks in low pressure environments. No improvements were found for the more massive particles shown in the sample, except for minor ones at lower pressures (Fig. 14(c)).
@@ 3.3 Tests and results: high-pressure sample
In Fig. 16 we show the relative momentum resolution and bias as a function of the true momentum for the three particle types present in the sample. At lower momenta ($p_{\text{true}}<1\text{GeV/}c$) the resolution is close to 2% for pions and muons while it is closer to 8% for the protons. In this momentum region the multiple scattering component of the resolution is dominant. This component is inversely proportional to the particles’ $\beta$ factor. With a given momentum, the proton, having a higher mass, will always have a smaller $\beta$ and thus a worse expected resolution. Furthermore, while muons and protons at lower energies have the chance to produce longer and even looping tracks inside the detector, protons will tend to loose their energy more quickly, again due to their masses. This is clearly shown in Fig. 17(b) where the average particle lengths as a function of their true momentum is shown. Somewhat significant biases are also shown at these lower momenta, especially for protons. For momenta $p_{\text{true}}>1\leavevmode\nobreak\ \text{GeV/}c$, the resolution is comparable for the three particle types and increases slowly with the particle momentum. At higher momenta the point resolution component is dominant, so a direct proportionality on the momentum is expected, with no distinction between the particle types. An inverse dependency on the lever arm and number of points and thus indirectly on the length is also expected, but as shown in Fig. 17(b), in this momentum range the average length of the track becomes roughly the same for all particle types, since muons and pions have momenta that are too high to produce looping tracks, as shown in Fig. 13(b).
@@
Figure 17: (a) Average $\beta$ as a function of the track length $l$ and (b) average track length as a function of the momentum $p$ for the HP sample.
In Fig. 18 we show the momentum resolution and bias as a function of the true track lengths $l$ for the three particle types. As could be predicted from Eq. 36, an inverse proportionality of the relative resolution on $l$ can be observed for all particle types. The worse performance observed for the protons can be explained by the average $\beta$ shown in Fig. 17(a), which is always smaller for more massive particles regardless of the length of the tracks. For longer tracks the hit component of the resolution is dominant and the difference in $\beta$ is not as impactful. A somewhat significant bias can be seen at lower lengths. Similar dependencies on lever arm $L_{\textrm{Arm}}$ and number of points in the track $N$, which can be treated as a proxy for $l$ in most cases, are shown in Figs. 21 and 22 in the Appendix.
Figure 18: Similar plots to Fig. 16. In this case the relative momentum resolution (a) and bias (b) are shown as a function of the true track length $l$ for the HP sample.
@@ 4 Conclusions
We introduced a Kalman Filter tailored for homogeneous gas TPCs, adapted from the one used in the ALICE experiment. An inherent limitation in the original ALICE approach arises from its suitability only for tracks describing a semi-circle at most in the $xy$ plane, perpendicular to the detector’s magnetic field. We discovered that this challenge can be addressed by applying a mirror rotation to the state vector when reaching the semi-circle’s boundaries. This adjustment is facilitated by introducing an $xy$ plane rotation during the Kalman Filter’s propagation step, converting the longitudinal $x$ coordinate into the radial distance from the rotation center. By implementing this technique, the new algorithm can effectively track trajectories of any length, including multiple circular paths within the detector (loopers). This enhancement has no precedent in the literature and significantly improves upon the ALICE Kalman Filter, making it highly promising for both the ALICE TPC [43] and the neutrino HPgTPCs like ND-GAr, which, due to the randomness in the production points of charged particles coupled with their relatively low energy, are prone to longer track formations. To evaluate the new algorithm, we developed a toy MC simulation tool named fastMCKalman and generated a sample with diverse detector and particle properties to validate its performance across a wide parameter space. Multiple tests conducted on this sample demonstrated that the algorithm’s estimates for parameter covariance effectively describe the sample and align closely with theoretical expectations. Furthermore, we examined the impact of the mirroring technique by comparing the ratios of $q/p_{\textrm{T}}$ resolutions with and without its application, revealing relative improvements by up to 80% for low-energy electrons and up to 50% for muons and pions. Additionally, we evaluated the new Kalman Filter algorithm’s performance using a sample of particles propagated in a high material budget environment, simulating conditions akin to a HPgTPC like ND-GAr. Realistic assessments of relative momentum resolution and bias showed behaviors consistent with theoretical expectations, affirming the viability of applying the method to a neutrino gas TPC.
@@ Acknowledgments
We define the reconstruction efficiency $\epsilon$ as the fraction of the correctly simulated tracks for which the CKF algorithm is fully propagated. In Figs. 19 and 20, we show the reconstruction efficiencies as a function of the initial true $p_{\textrm{T}}$, $L_{\textrm{Arm}}$ and $N$ for the PS and HP samples, respectively. The reconstruction efficiencies in both samples are shown to be very close to 100% except for very short low momentum tracks, for which other approaches than a Kalman Filter would most likely be used [25].
Figure 19: Portion of tracks in the PS sample for which the reconstruction was successful as a function of (a) initial $p_{\textrm{T}}$, (b) lever arm $L_{\textrm{Arm}}$ and (c) number of points in the track $N$.
Figure 20: Portion of tracks in the HP sample for which the reconstruction was successful as a function of (a) initial $p_{\textrm{T}}$, (b) lever arm $L_{\textrm{Arm}}$ and (c) number of points in the track $N$.
In Figs. 21 and 22, we show the relative momentum resolution and bias for the three particle types present in the HP sample, as a function of $L_{\textrm{Arm}}$ and $N$, respectively. These are analogous to Fig. 18 in the main text.
Figure 21: Similar plots to Fig. 16. In this case the relative momentum resolution (a) and bias (b) are shown as a function of the true track lever arm $L_{\textrm{Arm}}$ for the HP sample.
Figure 22: Similar plots to Fig. 16. In this case the relative momentum resolution (a) and bias (b) are shown as a function of the number of points in the track $N$ for the HP sample.

### v2

@@ 3.2 Tests and results: parameter scan sample
The PS sample was further used to test the improvement in $q/p_{\text{T}}$ resolution brought by the introduction of the “mirror rotation” technique in the CKF, compared to BKF. The fraction of the total tracks for which the mirroring technique was used $\epsilon_{\textrm{Mirror}}$ is shown in Fig. 13 as a function of initial true $p_{\textrm{T}}$ and $P_{\textrm{gas}}$. We show the fraction for the primaries in the first row and for the secondaries in the second. The particle types are divided in order of mass between electrons in Figs. 13(a) and 13(d), muons and pions in Figs. 13(b) and 13(e), kaons and protons in Figs. 13(c) and 13(f). We can see from the upper row plots and the low $p_{\textrm{T}}$ component of the lower row plots that the likelihood of the particles producing looping trajectory drops significantly with mass, due to the higher $\textrm{d}E/\textrm{d}x$. At low momenta the pressure of the gas also becomes important. This is especially true for the higher mass particles such as protons and kaons, which at higher $P_{\textrm{gas}}$ are stopped in the detector before producing any looping trajectory (see Fig. 13(c)). From Figs. 13(a), 13(b) and 13(c) it can also be shown that the only primary particle that necessitate the use of the mirroring operation are those that produce a looping trajectory in the detector, which is only possible at low initial transverse momenta $p_{\textrm{T}}<0.3\ \textrm{GeV}/c$. Secondary particle trajectories, on the other hand, can cover more than a semi-plane even if they don’t belong to loopers and can thus be produced at any momenta, as shown in Figs. 13(d), 13(e) and 13(f). At low transverse momenta, $\epsilon_{\textrm{Mirror}}$ is still significantly higher for secondaries.
The reconstruction efficiency $\epsilon$, defined as the fraction of the correctly simulated tracks for which the algorithm is fully propagated, was tested for CKF and BKF. It is shown as a function of the initial true $p_{\textrm{T}}$ and the $N$ of the total track in the first and second column of Fig 14 respectively. In Figs. 14(a) and 14(b) only the tracks for which the mirroring technique is needed to reconstruct all the points are shown, while the other tracks are shown Figs. 14(c) and 14(d). For the tracks that can be fully reconstructed without using the mirroring technique, the efficiency is essentially identical between the BKF and CKF algorithm. It is shown to be very close to 1 except for very low $p_{\textrm{T}}$ and $N$, for which other approaches than a Kalman Filter would most likely be used [25]. On the other hand $\epsilon$ is significantly improved by the CKF for low $N$ “mirrored” tracks, in some cases going from an efficiency of $\sim 0.5$ to $\epsilon>0.9$. This is a direct consequence of the increased number of points becoming available through the use of the mirroring technique.
Figure 13: Portion of tracks in the PS sample for which the mirror rotation was applied, $\epsilon_{\textrm{Mirror}}$, as a function of the gas pressure $P_{\textrm{gas}}$ and the initial true transverse momentum $p_{\textrm{T}}$ of the particle. The primaries are shown in the upper row, while the secondaries are shown in the lower row. The particle types are separated based on their mass: (a) and (d) contain only electrons, (b) and (e) muons and pions, (c) and (f) kaons and protons.
Figure 14: Reconstruction efficiency $\epsilon$ for the PS sample, as a function of the total number of points in the track $N$ and the initial true transverse momentum $p_{\textrm{T}}$. The results for the tracks for which the mirroring technique is needed to reconstruct all the points, are shown in (a) and (b) for the CKF and BKF respectively. The results for the other tracks are shown in (c) for the CKF and (d) for the BKF.
@@
Figure 15: Ratios of the $q/p_{\text{T}}$ resolutions obtained using the full CKF algorithm including the mirror rotation method $\sqrt{C_{44}^{\textrm{CKF}}}$, over the reconstruction without mirror rotation $\sqrt{C_{44}^{\text{BKF}}}$. The plots were produced for the whole PS sample. The histograms are color coded according to the gas pressure $P_{\textrm{gas}}$. Plot (a) contains only electrons, plot (b) contains pions and muons and plot (c) contains kaons and protons. Only tracks with a minimum of 30 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
Figure 16: Ratios of the $q/p_{\text{T}}$ resolutions obtained using the full CKF algorithm including the mirror rotation method $\sqrt{C_{44}^{\textrm{CKF}}}$, over the reconstruction without mirror rotation $\sqrt{C_{44}^{\text{BKF}}}$. The plots were produced for the whole PS sample. The histograms are color coded according to number of points in the tracks $N$. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
The difference between the results is shown to be less dramatic in more pressurized environments, where particles tend to be absorbed sooner and tracks are generally shorter. This trend is confirmed looking at the results obtained for the muons and pions sample in Fig. 15(b), for which $\textrm{d}E/\textrm{d}x$ will on average be higher due to their higher masses: the improvement in this case is by $\sim 20\%$ for $p_{\text{T}}<150$ MeV/$c$ with peaks of up to $\sim 50$% for the lowest momentum tracks in low pressure environments. No improvements were found for the more massive particles shown in the sample, except for minor ones at lower pressures (Fig. 15(c)).
@@ 3.3 Tests and results: high-pressure sample
The reconstruction efficiency $\epsilon$ was also tested. The results are shown in the Appendix in Fig. 20 and are similar to the PS Sample.
@@
In Fig. 17 we show the relative momentum resolution and bias as a function of the true momentum for the three particle types present in the sample. At lower momenta ($p_{\text{true}}<1\text{GeV/}c$) the resolution is close to 2% for pions and muons while it is closer to 8% for the protons. In this momentum region the multiple scattering component of the resolution is dominant. This component is inversely proportional to the particles’ $\beta$ factor. With a given momentum, the proton, having a higher mass, will always have a smaller $\beta$ and thus a worse expected resolution. Furthermore, while muons and protons at lower energies have the chance to produce longer and even looping tracks inside the detector, protons will tend to loose their energy more quickly, again due to their masses. This is clearly shown in Fig. 18(b) where the average particle lengths as a function of their true momentum is shown. Somewhat significant biases are also shown at these lower momenta, especially for protons. For momenta $p_{\text{true}}>1\leavevmode\nobreak\ \text{GeV/}c$, the resolution is comparable for the three particle types and increases slowly with the particle momentum. At higher momenta the point resolution component is dominant, so a direct proportionality on the momentum is expected, with no distinction between the particle types. An inverse dependency on the lever arm and number of points and thus indirectly on the length is also expected, but as shown in Fig. 18(b), in this momentum range the average length of the track becomes roughly the same for all particle types.
@@
Figure 18: (a) Average $\beta$ as a function of the track length $l$ and (b) average track length as a function of the true momentum $p_{\textrm{true}}$ for the HP sample.
In Fig. 19 we show the momentum resolution and bias as a function of the true track lengths $l$ for the three particle types. As could be predicted from Eq. 36, an inverse proportionality of the relative resolution on $l$ can be observed for all particle types. The worse performance observed for the protons can be explained by the average $\beta$ shown in Fig. 18(a), which is always smaller for more massive particles regardless of the length of the tracks. For longer tracks the hit component of the resolution is dominant and the difference in $\beta$ is not as impactful. A somewhat significant bias can be seen at lower lengths. Similar dependencies on lever arm $L_{\textrm{Arm}}$ and number of points in the track $N$, which can be treated as a proxy for $l$ in most cases, are shown in Figs. 22 and 23 in the Appendix. Similar plots have also been produced for the PS sample to allow for direct comparison in Figs. 24 and 25 also in the Appendix.
Figure 19: Relative momentum resolution (a) and bias (b) as function of the true track length, $l$, for the HP sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The three particle types (protons, muons and pions) are drawn separately.
@@ 4 Conclusions
In this paper we introduced a Kalman Filter tailored for homogeneous gas TPCs, adapted from the one used in the ALICE experiment. An inherent limitation in the original ALICE approach arises from its suitability only for tracks describing a semi-circle at most in the $xy$ plane, perpendicular to the detector’s magnetic field. We discovered that this challenge can be addressed by applying a mirror rotation to the state vector when reaching the semi-circle’s boundaries. This adjustment is facilitated by introducing an $xy$ plane rotation during the Kalman Filter’s propagation step, converting the longitudinal $x$ coordinate into the radial distance from the rotation center. By implementing this technique, the new algorithm can effectively track trajectories of any length, including multiple circular paths within the detector (loopers). This enhancement has no precedent in the literature and significantly improves upon the ALICE Kalman Filter, making it highly promising for both the ALICE TPC [47] and neutrino HPgTPCs like ND-GAr, which, due to the randomness in the production points of charged particles coupled with their relatively low energy, are prone to longer track formations. To evaluate the new algorithm, we developed a toy MC simulation tool named fastMCKalman and generated a sample with diverse detector and particle properties to validate its performance across a wide parameter space. Multiple tests conducted on this sample demonstrated that the algorithm’s estimates for parameter covariance effectively describe the sample and align closely with theoretical expectations.This last point is highlighted by Figs. 10, 11 and 12 where the theoretically expected resolution was first shown on its own for different particle types, gas densities and resolution and then directly compared with the CKF estimations by showing their ratios. Despite the wide resolution ranges shown for the theoretical expectations, the ratios are uniformly close to 1, demonstrating close agreement. A significant improvement in the reconstruction efficiency for the low $N$ and low $p_{\textrm{T}}$ mirrored tracks was also shown, in some cases going from $\epsilon\sim 0.5$ to $\epsilon>0.9$. Furthermore, we examined the impact of the mirroring technique by comparing the ratios of $q/p_{\textrm{T}}$ resolutions with and without its application, revealing relative improvements by up to 80% for low-energy electrons and up to 50% for muons and pions. Additionally, we evaluated the new Kalman Filter algorithm’s performance using a sample of particles propagated in a high material budget environment, simulating conditions akin to a HPgTPC like ND-GAr. Realistic assessments of relative momentum resolution and bias showed behaviors consistent with theoretical expectations, affirming the viability of applying the method to a neutrino gas TPC.
The reconstruction efficiency and resolution improvements brought by the application of the new algorithm to such a detector are very attractive. The improvement in $\mu$ and $\pi$ reconstruction would be extremely valuable when studying nuclear effects using techniques such as TKI, the effectiveness of which is directly dependent on the momentum reconstruction quality. The significant improvement seen for low-momentum electrons will be crucial for DUNE as it allows to more efficiently probe the $\nu_{e}$ and $\bar{\nu}_{e}$ fluxes. Additionally, it is likely that the mirroring technique could be applied to track formation. In most track forming algorithms, looping tracks are formed in segments, which are then connected together using various positional arguments. However, this step produces additional combinatorial complexity and makes algorithms more computationally heavy. The mirroring technique could be used to produce looping tracks in a more seamless way, expanding the Kalman Filter approach, which involves adding points one by one using track projection and compatibility $\chi^{2}$ arguments. This is for example the method currently used by the ALICE experiment [26, 27], which could directly benefit from this expansion.
@@ Acknowledgments
The reconstruction efficiency for the HP sample is shown in Fig. 20. Analogously to what was shown in Fig. 14, $\epsilon$ is plotted as a function of the initial true $p_{\textrm{T}}$ and the $N$. The results for the CKF and BKF are shown in the first and second column respectively. In the plots occupying the first row only the tracks for which the mirroring technique is used are shown, while the other tracks are shown in the plots in the second row. The results are analogous to what was shown for the PS sample in Fig. 14.
Figure 20: Reconstruction efficiency $\epsilon$ for the HP sample, as a function of the total number of points in the track $N$ and the initial true transverse momentum $p_{\textrm{T}}$. Analogous to Fig. 14
In Fig. 21 we show an alternative version of the plot in Fig.10 which focuses on the more statistically rich central region.
In Figs. 22 and 23, we show the relative momentum resolution and bias for the three particle types present in the HP sample, as a function of $L_{\textrm{Arm}}$ and $N$, respectively. These are analogous to Fig. 19 in the main text. Similar plots are produced for the PS sample to allow for a direct comparison. In Figs. 24 and 25, we show the relative momentum resolution and bias for the particles in the PS sample as a function of $p$ and $l$ respectively. The particle types are divided by mass into protons and kaons, pions and muons and finally electrons. Note that in this case the resolution and material properties are not uniform in the sample.
Figure 21: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to their particle type. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
Figure 22: Relative momentum resolution (a) and bias (b) as function of of the tracks’ lever arm $L_{\textrm{Arm}}$ for the HP sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The three particle types (protons, muons and pions) are drawn separately.
Figure 23: Relative momentum resolution (a) and bias (b) as function of as a function of the number of points in the tracks $N$ for the HP sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The three particle types (protons, muons and pions) are drawn separately.
Figure 24: Relative momentum resolution (a) and bias (b) as function of the true momentum, $p_{\textrm{true}}$, for the PS sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The particle types are divided based on their mass (protons and kaons, muons and pions, electrons) are drawn separately.
Figure 25: Relative momentum resolution (a) and bias (b) as function of the track length, $l$, for the PS sample. The two properties are defined as $\mu$ and $\sigma$ of Gaussian fits of the momentum fractional residuals $p_{\text{reco}}/p_{\text{true}}-1$. The particle types are divided based on their mass (protons and kaons, muons and pions, electrons) are drawn separately.